# OpenDebateEvidence: A Massive-Scale Argument Mining and Summarization Dataset

### v2

@@ 1 Introduction
  5. 5.
    We discuss the practical applications of OpenDebateEvidence in various fields such as legal document analysis, educational tools, and AI model development, highlighting its real-world relevance and impact.
@@ 3.1 Data Collection
OpenDebateEvidence is sourced from the OpenCaseList project (Hardy, 2024), an online platform where high school and college debate teams disclose and open-source their evidence. The dataset contains over 3.5 million documents, covering all NSDA debate topics from 2012 to 2023.333A list of these topics can be found here. Each document corresponds to a single piece of evidence used in a debate, categorized by debate format (Policy, LD, Public Forum), and includes comprehensive metadata such as author, date, title, source, citation details, and the debate round in which it was used [^4] An example of some of these downloads can be found here.
@@ 3.4 Data Statistics
The OpenDebateEvidence dataset offers a comprehensive collection of over 3.5 million documents categorized by debate format (Policy, Lincoln-Douglas, and Public Forum). Each document is enriched with extensive metadata, including author, date, title, source, citation details, and debate round information. Standardized tags describe the type of argument, such as topicality, disadvantages, advantages, and counterplans.
Policy Debate evidence constitutes approximately two-thirds of the dataset, Lincoln-Douglas Debate evidence comprises about one-third, and Public Forum Debate evidence makes up a smaller percentage. Spanning topics from 2012 to 2023, the dataset represents over 1,400 schools and includes contributions from more than 3,200 authors.
@@
Total Documents 3,571,098
Policy Debate Evidence 2,380,600
Lincoln-Douglas Debate Evidence 1,164,132
Public Forum Debate Evidence 26,366
Years Covered 2012-2023
Average Document Length (characters) 3,556
Total Schools Represented 1,423
Unique Authors 3,217
@@ 4.1.2 LLM as Judge
4.2 Results
The results of our experiments are shown in Tables 2, 3 and 4. Both the Mistral-7B and LLaMA3-8B models achieved promising results, with the LLaMA3-8B models generally outperforming the Mistral-7B models across all ROUGE metrics. Notably, fine-tuning on a larger subset of the dataset ($1{,}000{,}000$ examples) significantly improved performance, as evidenced by the LLaMA3-8B LoRA (1M Ex) model. This model achieved the highest scores across traditional NLP metrics and the LLM as Judge evaluations, demonstrating that extensive fine-tuning on domain-specific data is crucial for optimizing model performance in argument mining tasks.
Our results demonstrate the significant impact of using the OpenDebateEvidence dataset on model performance. On all three datasets, our fine-tuned models showed substantial improvements over the baseline models. The LLaMA3-8B LoRA model, fine-tuned on $1{,}000{,}000$ examples from OpenDebateEvidence, achieved the highest scores across all traditional NLP metrics and the LLM as Judge evaluations. This highlights the importance of extensive fine-tuning on domain-specific data.
4.2.1 OpenDebateEvidence Performance
For the OpenDebateEvidence dataset (Table 2), the LLaMA3-8B models generally outperformed the Mistral-7B models across all ROUGE metrics. Fine-tuning techniques, particularly LoRA with $1{,}000{,}000$ examples, significantly enhanced the models’ ability to generate high-quality summaries that effectively support arguments. Notably, the LoRA technique improved performance by effectively adapting model parameters with minimal additional computational resources. ReFT showed strong performance, indicating its ability to modify hidden representations in targeted subspaces, improving summarization quality. Orthogonalization, while effective, showed relatively less improvement than LoRA and ReFT, likely due to its focus on controlling specific features in the residual stream.
@@ 4.2.1 OpenDebateEvidence Performance
Model R-1 R-2 R-L Perplexity Output Quality Support Quality
@@
LoRA $30.1\pm 0.5$ $9.4\pm 0.2$ $25.8\pm 0.6$ $33.9\pm 2.3$ $7.7\pm 0.2$ $7.5\pm 0.2$
@@
4.2.2 BillSum Performance
Similarly, for the BillSum dataset (Table 3), the models fine-tuned on OpenDebateEvidence demonstrated superior performance compared to their baselines. This suggests that training on OpenDebateEvidence can enhance the models’ capabilities in reasoning and summarization, even when applied to a different domain such as legislative texts. The results on BillSum confirm the transferability of the fine-tuning techniques, with LoRA once again showing the most significant improvements. This technique’s efficiency in parameter adaptation appears to be particularly beneficial for handling diverse datasets. ReFT also performed well, indicating its robustness in capturing complex argument structures across different domains. Orthogonalization, while showing improvement, was less impactful compared to the other techniques.
@@ 4.2.2 BillSum Performance
Model R-1 R-2 R-L Perplexity Output Quality Support Quality
@@
LoRA $47.1\pm 0.5$ $23.4\pm 0.2$ $42.8\pm 0.6$ $23.9\pm 0.3$ $7.4\pm 0.2$ $7.2\pm 0.2$
ReFT $46.9\pm 0.2$ $23.3\pm 0.1$ $42.6\pm 0.6$ $24.3\pm 0.4$ $7.5\pm 0.3$ $7.3\pm 0.3$
@@
4.2.3 DebateSum Performance
In evaluating the DebateSum dataset (Table 4), the fine-tuned models demonstrated notable improvements over their base counterparts. The use of advanced fine-tuning techniques, particularly LoRA with a larger dataset of $1{,}000{,}000$ examples, significantly boosted the models’ performance across all ROUGE metrics. This was evident in both the Mistral-7B and LLaMA3-8B models, highlighting the effectiveness of LoRA in enhancing summarization capabilities. The ReFT technique also showed robust results, suggesting its strong ability to refine hidden representations for better summarization quality. While Orthogonalization improved the performance, it was comparatively less effective than LoRA and ReFT, likely due to its narrower focus on specific feature control within the residual streams.
@@ 4.2.3 DebateSum Performance
Model R-1 R-2 R-L Perplexity Output Quality Support Quality
@@
LoRA $28.5\pm 0.6$ $8.9\pm 0.3$ $24.7\pm 0.5$ $30.2\pm 2.0$ $7.5\pm 0.3$ $7.3\pm 0.3$
@@
5 Potential Applications and Future Directions
5.1 Argument Quality Assessment
The extensive collection of real-world debate arguments and rich metadata in the dataset provides a unique opportunity to study argument quality. By analyzing factors such as duplicate count, position in the debate round, and debate outcomes, researchers can develop models to automatically assess persuasiveness, relevance, and overall quality. This could lead to tools that offer real-time feedback and suggestions for debaters, improving debate communities’ inclusivity by reducing reliance on expensive human judges.
5.2 Multi-level Argument Summarization
The hierarchical structure of arguments in the dataset enables research into multi-level argument summarization. Models can generate summaries at various granularities, from concise one-sentence summaries to detailed overviews. This aligns with the emerging interest in query-focused and hierarchical summarization in the NLP community.
5.3 Argument Generation and Rebuttal
With its diverse collection of arguments and counterarguments, OpenDebateEvidence is valuable for developing argument generation models. By studying successful debaters’ patterns and strategies, researchers can create systems that generate persuasive and relevant arguments on given topics. Additionally, the dataset’s balanced coverage of affirmative and negative arguments enables the development of rebuttal generation models that counter opposing arguments.
5.4 Cross-domain Argument Mining
While primarily focused on competitive debate formats, the argumentation skills and techniques in the dataset are applicable across domains such as legal reasoning, policy-making, and online discussions. Researchers can develop general argument mining models for diverse argumentative texts, advancing areas like legal document analysis, opinion mining, and fact-checking. Integrating OpenDebateEvidence with fact-checking and misinformation detection datasets could yield robust models for identifying and countering misleading claims in public discourse.
5.5 Understanding Persuasion and Sentiment
The dataset captures both arguments’ logical structure and debaters’ rhetorical strategies and emotional appeals. By studying the interplay between rational argumentation and affective language, researchers can develop sophisticated models for understanding sentiment’s role in persuasion and decision-making. This has applications in political science, marketing, and human-computer interaction.
5.6 Debate Coaching and Education
OpenDebateEvidence holds significant potential for debate coaching and education. Analyzing successful arguments and strategies can help coaches identify best practices and develop more effective training programs. The dataset can also serve as a resource for creating educational materials such as argument templates, case studies, and interactive learning tools. This will support aspiring debaters in skill development.

### v3

@@ 3.1 Data Collection
OpenDebateEvidence is sourced from the OpenCaseList project (Hardy, 2024), an online platform where high school and college debate teams disclose and open-source their evidence. The dataset contains over 3.5 million documents, covering all NSDA debate topics from 2014 to 2022.333A list of these topics can be found here. Each document corresponds to a single piece of evidence used in a debate, categorized by debate format (Policy, LD, Public Forum), and includes comprehensive metadata such as author, date, title, source, citation details, and the debate round in which it was used [^4] An example of some of these downloads can be found here.
@@ 3.3 Data Deduplication
We note that this is a description of how we created the "Duplicates" metadata column. We performed a separate neural semantic de-duplication procedure which is described in the Appendix.
@@ 3.4 Data Statistics
The OpenDebateEvidence dataset offers a comprehensive collection of over 3.5 million documents categorized by debate format (Policy, Lincoln-Douglas, and Public Forum). Each document is enriched with over 40 columns of extensive metadata, including author, date, title, source, citation details, and debate round information. Standardized tags describe the type of argument, such as topicality, disadvantages, advantages, and counterplans.
Policy Debate evidence constitutes approximately two-thirds of the dataset, Lincoln-Douglas Debate evidence comprises about one-third, and Public Forum Debate evidence makes up a smaller percentage. Spanning topics from 2014 to 2022, the dataset represents over 1,300 schools and includes contributions from more than 6,400 authors.
@@
Total Rows 4,830,561
Total Documents (valid full-text column) 3,512,280
Policy Debate Evidence 2,768,419
Lincoln-Douglas Debate Evidence 1,526,383
Public Forum Debate Evidence 43,131
Years Covered 2014-2022
Average Document Length (characters) 5,270
Total Schools Represented 1,366
Unique Authors 6,455
@@ 4.1.2 LLM as Judge
Model R-1 (%) R-2 (%) R-L (%) Perplexity Output Quality Support Quality
@@
LoRA $\mathbf{30.1\pm 0.5}$ $\mathbf{9.4\pm 0.2}$ $\mathbf{25.8\pm 0.6}$ $\mathbf{33.9\pm 2.3}$ $\mathbf{7.7\pm 0.2}$ $\mathbf{7.5\pm 0.2}$
@@
LLaMA3-70B
Base $33.8\pm 1.2$ $14.1\pm 0.9$ $30.2\pm 1.3$ $45.7\pm 3.2$ $7.9\pm 0.2$ $7.5\pm 0.2$
LoRA $\mathbf{37.2\pm 1.0}$ $\mathbf{15.8\pm 0.8}$ $\mathbf{33.4\pm 1.1}$ $\mathbf{27.3\pm 2.7}$ $\mathbf{8.2\pm 0.2}$ $\mathbf{8.1\pm 0.2}$
ReFT $35.9\pm 1.1$ $15.1\pm 0.7$ $32.6\pm 1.2$ $31.4\pm 2.8$ $8.0\pm 0.3$ $7.9\pm 0.2$
Orthogonal $34.2\pm 1.2$ $14.3\pm 0.8$ $31.1\pm 1.3$ $39.9\pm 3.1$ $8.1\pm 0.2$ $7.7\pm 0.3$
Google Gemini
Base $32.5\pm 1.3$ $13.5\pm 0.9$ $29.4\pm 1.4$ $49.8\pm 3.6$ $\mathbf{8.5\pm 0.2}$ $7.9\pm 0.3$
Anthropic Claude
Base $31.2\pm 1.4$ $12.9\pm 0.9$ $28.1\pm 1.3$ $52.1\pm 3.7$ $\mathbf{8.7\pm 0.3}$ $7.8\pm 0.3$
@@
Model R-1 (%) R-2 (%) R-L (%) Perplexity Output Quality Support Quality
@@
LoRA $\mathbf{47.1\pm 0.5}$ $\mathbf{23.4\pm 0.2}$ $\mathbf{42.8\pm 0.6}$ $\mathbf{23.9\pm 0.3}$ $7.4\pm 0.2$ $7.2\pm 0.2$
ReFT $46.9\pm 0.2$ $23.3\pm 0.1$ $42.6\pm 0.6$ $24.3\pm 0.4$ $\mathbf{7.5\pm 0.3}$ $\mathbf{7.3\pm 0.3}$
@@
LLaMA3-70B
Base $50.2\pm 1.1$ $27.4\pm 0.9$ $45.7\pm 1.2$ $22.9\pm 2.3$ $7.8\pm 0.2$ $7.6\pm 0.2$
LoRA $\mathbf{54.6\pm 1.0}$ $\mathbf{30.5\pm 0.8}$ $\mathbf{50.0\pm 1.1}$ $\mathbf{19.7\pm 2.1}$ $\mathbf{8.0\pm 0.2}$ $\mathbf{8.0\pm 0.2}$
ReFT $52.9\pm 1.1$ $29.1\pm 0.7$ $48.3\pm 1.2$ $20.5\pm 2.4$ $7.9\pm 0.3$ $7.8\pm 0.2$
Orthogonal $51.1\pm 1.2$ $28.0\pm 0.9$ $46.6\pm 1.4$ $23.8\pm 2.6$ $7.9\pm 0.3$ $7.7\pm 0.3$
Google Gemini
Base $48.7\pm 1.3$ $26.0\pm 1.0$ $44.3\pm 1.5$ $24.9\pm 2.8$ $\mathbf{8.4\pm 0.3}$ $\mathbf{8.0\pm 0.3}$
Anthropic Claude
Base $47.3\pm 1.4$ $24.8\pm 1.1$ $42.9\pm 1.6$ $26.7\pm 3.0$ $\mathbf{8.3\pm 0.3}$ $7.8\pm 0.3$
@@
Model R-1 (%) R-2 (%) R-L (%) Perplexity Output Quality Support Quality
@@
LoRA $\mathbf{28.5\pm 0.6}$ $\mathbf{8.9\pm 0.3}$ $\mathbf{24.7\pm 0.5}$ $\mathbf{30.2\pm 2.0}$ $\mathbf{7.5\pm 0.3}$ $\mathbf{7.3\pm 0.3}$
@@
LLaMA3-70B
Base $36.7\pm 1.2$ $16.9\pm 0.9$ $32.8\pm 1.3$ $42.1\pm 3.1$ $7.7\pm 0.2$ $7.4\pm 0.2$
LoRA $\mathbf{39.4\pm 1.1}$ $\mathbf{18.5\pm 0.8}$ $\mathbf{35.1\pm 1.2}$ $\mathbf{28.9\pm 2.5}$ $\mathbf{8.0\pm 0.2}$ $\mathbf{8.0\pm 0.2}$
ReFT $38.1\pm 1.2$ $17.8\pm 0.7$ $34.2\pm 1.2$ $30.5\pm 2.8$ $7.9\pm 0.3$ $7.8\pm 0.2$
Orthogonal $36.9\pm 1.3$ $17.0\pm 0.8$ $33.0\pm 1.4$ $39.3\pm 3.2$ $7.8\pm 0.3$ $7.6\pm 0.3$
Google Gemini
Base $35.5\pm 1.4$ $15.8\pm 0.9$ $31.4\pm 1.5$ $44.5\pm 3.4$ $\mathbf{8.5\pm 0.3}$ $7.9\pm 0.3$
Anthropic Claude
Base $34.2\pm 1.5$ $15.0\pm 1.0$ $30.1\pm 1.6$ $46.8\pm 3.6$ $\mathbf{8.8\pm 0.3}$ $7.8\pm 0.3$
4.1.3 OpenDebateEvidence Performance
For the OpenDebateEvidence dataset (Table 2), the LLaMA3-70B models significantly outperformed both the LLaMA3-8B and Mistral-7B models across all ROUGE metrics. The larger model capacity of LLaMA3-70B contributed to its superior performance, highlighting the importance of model size in complex summarization tasks. Fine-tuning techniques, particularly LoRA, further enhanced the LLaMA3-70B model’s performance, with the LoRA fine-tuned model achieving the highest scores across all metrics. This underscores LoRA’s effectiveness in adapting large models with minimal additional parameters, allowing for improved summarization quality without extensive computational resources. ReFT also showed strong performance improvements on the LLaMA3-70B model, indicating its robustness in refining hidden representations for better output. Orthogonalization, while still providing gains, was less impactful compared to LoRA and ReFT.
The base versions of Google Gemini and Anthropic Claude models also demonstrated competitive performance, outperforming the base LLaMA3-8B and Mistral-7B models. However, they did not surpass the fine-tuned LLaMA3-70B models, suggesting that while these models are strong out-of-the-box, fine-tuning large models like LLaMA3-70B with techniques such as LoRA can yield superior results.
In summary, the new experimental results demonstrate the effectiveness of fine-tuning techniques, especially LoRA, when applied to larger models like LLaMA3-70B. These techniques consistently improved performance across all datasets, highlighting their utility in enhancing large language models for complex summarization tasks. The base versions of Google Gemini and Anthropic Claude showed strong capabilities, but targeted fine-tuning of large models remains crucial for achieving state-of-the-art results.
4.2 Additional Experiments
We conducted additional experiments to further evaluate the robustness and generalizability of our fine-tuning techniques. These experiments included assessments on the xSum dataset, a non-argumentative summarization task, and an analysis of inter-human and human-GPT-4 agreement rates in evaluating summary quality. We also conducted further experiments to evaluate our models on argument detection, classification tasks, stance detection, and downstream tasks. We also assessed the reliability of GPT-4 as an evaluator by comparing its ratings with those of human experts. These additional experiments are included in the appendix